[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun May 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon May 8 00:27:52 UTC 2017


HSU Shift Log for Sun May 07 2017 20:27:46 GMT-0400 (EDT)

Shift: R

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues: Got the red "guidance not ready" banner for EP90 at 00Z.  
Retrieved guidance and did not get SHIPS/LGEM the first time. Submitted  
guidance again at 0021Z and it was ready after 76 seconds. The 18Z HWRF for  
EP90 only returned a 0 hour tracker, so no HWFI was available for 00Z. I  
noticed that all of the SHIPS RII values at 18Z and 00Z for all thresholds  
were 0%, which seemed odd. The "%Contribution" value for all but one of the  
predictors was 999.0, which might indicate something.

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170508/2df35999/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list